A father is four times his son's age.And in 30yrs the son will be half as old as his father.How old are the father and son each now?
Answer by mananth(16946)

let son's age be x
the father's age = 4x
after 30 years
son's age = x+30
father's age = 4x+30
1/2* 4x+30 = x+30
4x+30 = 2*(x+30)
4x+30= 2x+60
4x-2x= 60-30
2x=30
x=15 son's age
father's age = 4x= 4*15 = 60 years